OpenStudy (anonymous):

Given two rectangles, one with length = l and width = w, and the second rectangle with length = 2l and width = 4w. How much larger will the perimeter of the second rectangle be than the perimeter of the first rectangle?Select one of the options below as your answer: A. P2 = P1 + 4w B. P2 = 2P1 + w C. P2 = 2P1 + 4w D. P2 = P1 + w

OpenStudy (kira_yamato):

P(1) = 2 (l + w) P(2) = 2 (2l + 4w) = 4 (l + w + w) = 4 (l + w) + 4w = 2 P(1) + 4w (C)

OpenStudy (anonymous):

answer is C for sure...!!